OpenStudy (anonymous):

What is the standard form equation of the line passing through (1, –5) with a slope of –1?

OpenStudy (bahrom7893):

y+5=-1(x-1)

OpenStudy (bahrom7893):

get this into a standard form: y+5=-x+1 y+x=-4